OpenStudy (anonymous):

draw a graph that correctly corresponds to 3x + 9y = -18

OpenStudy (anonymous):

change it to "y=mx+b" (slope-intercept) form \[3x+9y=-18~~~\implies~~~y=mx+b\]\[:~~y={-1\over3}~x-2\]